# TrafficGamer: Reliable and Flexible Traffic Simulation for Safety-Critical Scenarios with Game-Theoretic Oracles

### v1

@@ Methods
  • $\Omega_{i}$ and $\mathcal{A}_{i}$ denote the spaces of observations and actions for a specific agent $i$. The observations include the position, velocity, heading, and partial map features in a neighbor region of the agent $i$. The actions consist of relative heading and acceleration, as described in [11].
@@
  • $r_{i}:\{\Omega_{i}\times\mathcal{A}_{i}\}_{i=1}^{{I}}\rightarrow\mathbb{R}$ denotes the agent-specific reward function that maps actions and observations from all agents to the reward of $i^{th}$ agent. We consider reward factors such as collision avoidance, lane deviation, and reaching the destination.
  • $c_{i}:\{\Omega_{i}\times\mathcal{A}_{i}\}_{i=1}^{{I}}\rightarrow\mathbb{R}$ denotes the agent-specific cost function that maps actions and observations from all agents to the cost of $i^{th}$ agent. We adapt vehicle distance as a constraint condition, additionally, the expectation of cumulative constraint functions $c_{i}$ must not exceed associated thresholds episodic constraint threshold $\delta$.
@@
  • $\gamma\in[0,1]$ and $p_{0}\in\Delta^{\mathcal{S}}$ denote the discount factor and the initial distribution.
@@
Figure 10: Presentation of the game type. In cooperative games, agents share a common interest and work together to optimize a global objective, demonstrating fully collaborative behavior. In contrast, zero-sum games are defined by strict competition, where one player’s gain is exactly equal to another’s loss, ensuring that the total rewards among all participants remain constant—typically zero. General-sum games, however, present a more complex scenario where the sum of all agents’ rewards can vary significantly. In these games, while agents may initially act competitively due to self-interest, they can also maximize their benefits by collaborating with others, blending competitive and cooperative dynamics.
In the framework of Dec-POCMDP, each agent is assigned an individual reward function and cost function, denoted as $r_{i}(\cdot)$ and $c_{i}(\cdot)$. This aligns a general-sum game structure [57], more complex and less explored than zero-sum and cooperative games [58, 59]. Figure 10 displays the differences between these three settings. Despite challenges, modeling a general-sum game better aligns with real-world driving scenarios since human drivers often prioritize their objects. These objects can be applied to satisfy conflicting interests among different AV agents. During the optimization process, human drivers’ behaviors inevitably affect others’ decisions. Accordingly, we consider the General Sum Markov Games (GS-MGs) under the Dec-POCMDP. For the agent $i$, the value function $V^{\boldsymbol{\pi},r}_{i,t}:\mathcal{S}\rightarrow\mathbb{R}$ , action-value function $Q^{\boldsymbol{\pi},r}_{i,t}:\mathcal{S}\times\mathcal{A}_{i}\rightarrow\mathbb{R}$ and advantage function $A^{\boldsymbol{\pi},r}_{i,t}:\mathcal{S}\times\mathcal{A}_{i}\rightarrow\mathbb{R}$ are represented by:
@@
where $-\boldsymbol{a}_{i}=\{\mathds{1}_{i^{\prime}\neq i}a_{i}^{\prime}\}_{i^{\prime}=1}^{{I}}$ [^2] Throughout this work, the bold symbols (e.g., $\boldsymbol{a}$) indicate a vector of variables while the unbold ones (e.g., $a$) represent a single variable. denotes the joint action performed by ${I}-1$ players (without $i$’th player) and $\boldsymbol{\pi}=\{\pi_{i}\}_{i=1}^{{I}}$ denotes the product policy.
@@
Unlike previous GS-MG solvers [60, 39, 40, 61, 57] that rely on an interactive environment, traffic simulation presents additional challenges since our algorithm can only utilize an offline database that records the behaviors of multiple drivers on open roads. Additionally, we incorporate constraints into the optimization process to learn the agents’ behavior under traffic rules. This problem can be formulated as Offline Multi-agent Constrained Reinforcement Learning (Offline MA-CRL). Using the offline database that records the behaviors of multiple drivers on open roads, our algorithm aims to learn the CCE policies of multiple AVs under various constraints that accurately reflect the human drivers’ behaviors in the dataset. Specifically, the problem can be summarized as follows:
@@
(Offline MA-CRL in GS-MGs.) let $\mathcal{D}_{o}=\{\mathcal{M}_{n},\tau_{n,1},\dots,\tau_{n,{I}}\}_{n=1}^{{N}}$ defines the offline dataset, where $n=[{N}]$ defines the number of scenario, $\mathcal{M}_{n}$ presents the game context in the $n^{th}$ scenario, $\boldsymbol{c}=\{{c_{0},c_{1},...,c_{i}}\}_{i=1}^{{I}}$ presents constraints and $\tau_{n,i}=\{o_{i,0},a_{i,0},\dots,o_{i,T},a_{i,T}\}$ denotes the trajectory of $i^{th}$ agent in the $n^{th}$ scenario. Given $\mathcal{D}_{o}$, the goal of our algorithm is to learn a $\boldsymbol{\hat{\pi_{\boldsymbol{c}}}}$ that satisfies the constraints with the following properties: 1) Exploitability: $\boldsymbol{\hat{\pi_{\boldsymbol{c}}}}$ satisfies the $\epsilon$-approximate CCE in definition 1, and 2) Fidelity: $\boldsymbol{\hat{\pi_{\boldsymbol{c}}}}$ must be consistent with the real driver’s policies such that $D_{f}(\boldsymbol{\hat{\pi_{\boldsymbol{c}}}},\boldsymbol{\pi}^{o})\leq\xi$ where $D_{f}$ and $\xi$ denote the divergence metric and a threshold.
In this work, to solve the offline MA-CRL problem in GS-MGs under the Dec-POCMDP (definition 2), we consider a model-based MA-CRL approach that 1) trains a generative world model to acquire AV environment features in a data-driven way, 2) converges to CCEs based on the predicted environment dynamics and predefined action space and 3) adjusts the level of competition for capturing diverse degrees of CCEs.
@@
For computation efficiency, we follow the decoder-only architecture in GPT [62] and implement the world model as a Auto-regressive Trajectory Decoder. As a sequential prediction model, our world model maps the previous state (e.g., $s_{t-1}$) and the action of each agent (e.g., $a_{1,t},\dots,a_{i,t}$) into the next state $s_{t}$. Unlike QCNet [50], which generates trajectories for future vehicle motions over a fixed period, our world model adopts an auto-regressive approach, predicting vehicles’ step-wise motion based on its previous predictions. This approach enables modeling how past movements influence future decisions.
@@
Figure 11: Illustration of the decoder network module. We have $N$ attention layers to decode accurate trajectories. 1) Agent-to-Map Cross Attention incorporates the map information $z^{m}=\xi(\mathcal{M})$ ($\xi$ is a map encoder) including all kinds of lane markings and types of polygons into the most recent agents’ features such that e.g., $\boldsymbol{z}^{s}_{i,t}=\text{CrossAttn}(\tau_{1,t},\dots,\tau_{{I},t},z^{m})$. 2) Agent-to-Temporal Cross Attention that embeds the historical information into the agents’ trajectories $\tau_{1,t+1},\dots,\tau_{{I},t+1}$ where $\tau_{1,t+1}$ consists of $a_{1,t}$ and $s_{1,t}$, e.g., $z^{{W}}_{i,t}=\text{CrossAttn}(\{z^{s}_{i,t-1},\dots,z^{s}_{i,t-{W}}\},z^{s}_{i,t})$, where $H$ indicates the length of encoding historical steps. 3) Agent-to-Neighbor Cross Attention that embeds the spatial-temporal features of the surrounding agents such that $z^{N}_{i,t}=\text{CrossAttn}(z^{{W}}_{i,t},\{z^{{W}}_{j,t}\}_{j\in N_{i}})$ and 4) Self-Agent Attention considers the agent’s entire predicted trajectory at each time step, capturing dependencies between distant elements, e.g., $z^{A}_{i,t}=\text{SelfAttn}(\{z^{N}_{i,t}\}^{T}_{t=0})$. The final state $s_{t+1}$ is represented by concatenating $s_{t}=\{\boldsymbol{z}^{A}_{i,w}\}_{w=t-{W},i=0}^{t,{I}}$ with the predicted $\boldsymbol{z}^{A}_{1,t+1},\dots,\boldsymbol{z}^{A}_{{I},t+1}$.
@@
Under the context of Dec-POCMDP, when $t=0$, $s_{0}$ captures the static game map $\mathcal{M}$ and initial features of all agents. when ${t>0},s_{t}=\{\boldsymbol{z}^{A}_{i,w}\}_{w=t-{W},i=0}^{t,{I}}$ captures the spatial-temporal information of all agents under the game map $\mathcal{M}$ in the previous $t-1$ time steps, and $a_{1,t},\dots,a_{i,t}$ denotes the acceleration and heading of agents in the current time step. Under this setting, our decoder is implemented by 1) Agent-to-Map Cross Attention, 2) Agent-to-Temporal Cross Attention, 3) Agent-to-Neighbor Cross Attention and 4) Self-Agent Attention incorporates map information , historical information , the spatial-temporal features of the surrounding agents , and the features of the agent itself in the temporal dimension thereby mapping $s_{t}$ and $\boldsymbol{a}_{1,\dots,{I},t}$ to $s_{t+1}$. The details of implementation are illustrated in Figure 11. In addition, our world model also includes two other important modules:
@@
The actor model predicts the actions (acceleration and heading) based on the state $s_{t}=\{\boldsymbol{z}^{s}_{i,w}\}_{w=t-{W},i=0}^{t,{I}}$ to satisfy latent traffic rules in the realistic driving scenarios. The actor model $\pi_{i,t}(a_{i,t}|o_{i,t})$ denotes the probability of the $i$’s agent generates an action $a_{t}$ such that ${\max\limits_{k}}\enspace\omega_{i}^{k}p(a_{t}|\mu_{i,t}^{k},b_{i,t}^{k})=\pi_{i,t}(a_{i,t}|o_{i,t})$ where 1) $\omega_{i}^{k}$ denotes a learnable coefficient and 2) $p_{k}$ denotes the $k$-th mixture component’s Laplace density. We constrain the output actions within a reasonable range to prevent irrational driving behaviors such as sudden acceleration or deceleration, sharp turns, etc. In the subsequent RL fine-tuning stage, we sample the $i$-th agent’s future trajectory as a weighted mixture of Laplace distributions by following [50, 63]:
@@
where $\omega_{i}^{k}$ can effectively act as the weighting coefficients and $\mu_{i,t}^{k}$ and $b_{i,t}^{k}$ characterize the mean position and the level of uncertainty of the $i$-th agent at the time step $t$.
@@
For each agent, our observation model maps $s_{t}$ and $a_{i,t}$ into agent-specific observations $\{\boldsymbol{o}_{i,t}\}_{i}^{I}$. The observation model is implemented by $o_{i,t}=f^{\text{MLP}}_{i}(z^{A}_{i,t})$, where MLP is the Multilayer Perceptron. Besides, this module considers the historical actions and observation of all agents such that $\boldsymbol{h}=\{(o_{i,\iota},a_{i,\iota})\}_{\iota=0,i=1}^{w,{I}}$.
@@
As our environment is structured as a multi-player competitive game with rewards and costs specific to each agent, inspired by [40, 43], we consider a decentralized update of each agent’s policy where we fix the rest ${I}-1$ agents’ policy $\boldsymbol{\pi}_{-i}$ and train policy $\pi_{i}$ to get the best response of agent $i$. In this work, we update $\pi_{i}$ by iteratively optimizing the following objective:
@@
where $\beta_{i}(o)=\frac{c}{\rho_{i}(o)}$ serves as an exploration bonus to less visited state. $c$ is a hyper-parameter and $\rho_{i}(o)$ denotes the density of visited observation $o$ [64], representing the probability of $o$ occurrences at time $t$ [65]. Such an optimistic V-learning objective served as an extension of the CCE-V-Learning algorithm [40, 38], which had been proven to converge to CCE under discrete environments, and we extend this algorithm to solve continuous decision-making problems.
@@
We follow [43] and incorporate Bregman divergence $\mathcal{B}_{\psi}(\cdot,\cdot)$ with respect to the mirror map $\psi$ such that $\mathcal{B}_{\psi}(x,y)=\psi(x)-\psi(y)-\langle\nabla\psi(y),x-y\rangle$ into the objective with convergence guarantees. Recent studies [44, 43, 41, 42] justified that the mirror decent approaches can solve different kinds of games in multi-player settings. To derive a more intuitive objective, we implement the mirror map as the negative entropy such that $\psi(x)=\sum p(x)\log p(x)$, and the objective (12) becomes:
@@
However, both the transition function $\mathcal{T}$ and policy of other players $\boldsymbol{\pi}_{-i,t}$ are not subject to optimize in the objective (12), and thus recent studies [66, 67] often consider the discounted causal entropy [68] $\sum_{t=0}^{T}\gamma^{t}\mathcal{H}[\pi(a_{i,t}|o_{i,t})]$. Similarly, instead of utilizing the computationally intractable trajectory-level KL-divergence $\mathcal{D}_{kl}(\pi_{i}\|\pi^{j-1}_{i})$, we consider the time-wise causal KL-divergence $\sum_{t=0}^{T}\gamma^{t}\mathcal{D}_{kl}[\pi_{i,t}(\cdot)\|\pi^{j-1}_{i,t}(\cdot)]$, and by substituting it and the equation (11) into the objective (12), we have:
@@
To effectively simulate complex traffic scenarios that include various vehicle types—such as cars, buses, and trucks—and diverse driving styles, such as aggressive and conservative driving, it is crucial to tailor the behavior of each agent and regulate the level of competition in the scenarios created. By subjecting the AV control system to these varied scenarios, we can more thoroughly assess the system’s robustness. This comprehensive evaluation helps in developing trustworthy AV vehicles capable of performing reliably in realistic traffic conditions.
To accurately represent diverse levels of scenarios characterized by different CCEs, we incorporate the constrained and risk-sensitive policy optimization into the multi-agent traffic simulation system.
@@
To dynamically adjust the intensity of CCEs, we request the agents to AV agents varying levels of driving constraints, thereby modulating the severity and nature of the driving conditions. Specifically, we expand the objective (15) by formulating the trade-off between rewards and costs under a constrained policy optimization objective:
@@
This strategy explicitly manages the risk sensitivity of driving behaviors, thereby deriving risk-seeking or risk-averse policies for each AV agent. This approach effectively promotes either aggressive or conservative driving policies to enhance the realism and variability of the scenarios. To drive risk-sensitive and constraints-satisfying policies for multiple agents, inspired by [55], we develop a risk-sensitive constraint to capture the uncertainty induced by environmental dynamics and extend the objective 16 as follows:
@@
By following [71], we parameterize the distribution with $N$ supporting quantiles and update these function via quantile regression [56], which acts as an asymmetric squared loss in an interval [$-\kappa,\kappa$] around zero:

### v2

@@ Methods
  • $\Omega_{i}$ and $\mathcal{A}_{i}$ denote the spaces of observations and actions for a specific agent $i$. The observations include the position, velocity, heading, and partial map features in the surrounding neighbor region of the agent $i$. The actions consist of relative heading and acceleration, as described in [11].
@@
  • $r_{i}:\{\Omega_{i}\times\mathcal{A}_{i}\}_{i=1}^{{I}}\rightarrow\mathbb{R}$ denotes the agent-specific reward function that maps actions and observations from all agents to the reward of $i^{th}$ agent. We considered reward factors such as collision avoidance, lane deviation, and reaching the destination, etc.
  • $c_{i}:\{\Omega_{i}\times\mathcal{A}_{i}\}_{i=1}^{{I}}\rightarrow\mathbb{R}$ denotes the agent-specific cost function that maps actions and observations from all agents to the cost of $i^{th}$ agent. We adapt vehicle distance as a constraint condition, additionally, the expectation of cumulative constraint functions $c_{i}$ must not exceed associated the thresholds episodic constraint threshold $\delta$.
@@
  • $\gamma\in[0,1]$ and $p_{0}\in\Delta^{\mathcal{S}}$ denote the discount factor and the initial distribution, respectively.
@@
Figure 10: Presentation of the game type. In cooperative games, agents share a common interest and work together to optimize their ability to a global objective, demonstrating fully collaborative behavior. In contrast, zero-sum games are defined by strict competition, where one player’s gain is exactly equal to another’s loss, ensuring that the total rewards among all participants remain constant—typically zero. General-sum games, however, present a more complex scenario where the sum of all agents’ rewards can vary significantly. While agents may initially act competitively in these games due to self-interest, they can also maximize their benefits by collaborating with others, blending competitive and cooperative dynamics.
In the framework of Dec-POCMDP, each agent is assigned an individual reward function and cost function, denoted as $r_{i}(\cdot)$ and $c_{i}(\cdot)$. This aligns with a general-sum game structure [57], which is more complex and less explored than zero-sum and cooperative games [58, 59]. Figure 10 displays the differences between these three settings. Despite the challenges, modeling a general-sum game better aligns with real-world driving scenarios as human drivers often prioritize their objects. These objects can be applied to satisfy conflicting interests among different AV agents. During the optimization process, human drivers’ behaviors inevitably affect others’ decisions. Accordingly, we consider the General Sum Markov Games (GS-MGs) under the Dec-POCMDP. For the agent $i$, the value function $V^{\boldsymbol{\pi},r}_{i,t}:\mathcal{S}\rightarrow\mathbb{R}$ , action-value function $Q^{\boldsymbol{\pi},r}_{i,t}:\mathcal{S}\times\mathcal{A}_{i}\rightarrow\mathbb{R}$ and advantage function $A^{\boldsymbol{\pi},r}_{i,t}:\mathcal{S}\times\mathcal{A}_{i}\rightarrow\mathbb{R}$ are represented by:
@@
where $-\boldsymbol{a}_{i}=\{\mathds{1}_{i^{\prime}\neq i}a_{i}^{\prime}\}_{i^{\prime}=1}^{{I}}$ [^3] Throughout this work, the bold symbols (e.g., $\boldsymbol{a}$) indicate a vector of variables, while the non-bold ones (e.g., $a$) represent a single variable. denotes the joint action performed by ${I}-1$ players (without $i$’th player) and $\boldsymbol{\pi}=\{\pi_{i}\}_{i=1}^{{I}}$ denotes the product policy.
@@
Unlike previous GS-MG solvers [60, 39, 40, 61, 57] that rely on an interactive environment, traffic simulation presents additional challenges, as our algorithm can only utilize an offline database with records of the behaviors of multiple drivers on open roads. Additionally, we incorporate constraints into the optimization process to allow the algorithm to learn the agents’ behavior under traffic rules. This problem can be formulated as Offline Multi-agent Constrained Reinforcement Learning (Offline MA-CRL). Using the offline database with records of the behaviors of multiple drivers on open roads, our algorithm aims to learn the CCE policies of multiple AVs under various constraints that accurately reflect the human drivers’ behaviors in the dataset. Specifically, the problem can be summarized as follows:
@@
(Offline MA-CRL in GS-MGs.) let $\mathcal{D}_{o}=\{\mathcal{M}_{n},\tau_{n,1},\dots,\tau_{n,{I}}\}_{n=1}^{{N}}$ defines the offline dataset, where $n=[{N}]$ defines the number of scenarios, $\mathcal{M}_{n}$ represents the game context in the $n^{th}$ scenario, $\boldsymbol{c}=\{{c_{0},c_{1},...,c_{i}}\}_{i=1}^{{I}}$ represents constraints and $\tau_{n,i}=\{o_{i,0},a_{i,0},\dots,o_{i,T},a_{i,T}\}$ denotes the trajectory of $i^{th}$ agent in the $n^{th}$ scenario. Given $\mathcal{D}_{o}$, the goal of our algorithm is to learn a $\boldsymbol{\hat{\pi_{\boldsymbol{c}}}}$ that satisfies the constraints with the following properties: (1) Exploitability: $\boldsymbol{\hat{\pi_{\boldsymbol{c}}}}$ satisfies the $\epsilon$-approximate CCE in Definition 1, and (2) Fidelity: $\boldsymbol{\hat{\pi_{\boldsymbol{c}}}}$ must be consistent with the real driver’s policies such that $D_{f}(\boldsymbol{\hat{\pi_{\boldsymbol{c}}}},\boldsymbol{\pi}^{o})\leq\xi$ where $D_{f}$ and $\xi$ denote the divergence metric and a threshold, respectively.
In this work, to solve the offline MA-CRL problem in GS-MGs under the Dec-POCMDP (Definition 2), we consider a model-based MA-CRL approach that (1) trains a generative world model to acquire AV environment features in a data-driven way, (2) converges to CCEs based on the predicted environment dynamics and predefined action space and (3) adjusts the level of competition for capturing diverse degrees of CCEs.
@@
For computation efficiency, we follow the decoder-only architecture in GPT [62] and implement the world model as an Auto-regressive Trajectory Decoder. As a sequential prediction model, our world model maps the previous state (e.g., $s_{t-1}$) and the action of each agent (e.g., $a_{1,t},\dots,a_{i,t}$) onto the next state $s_{t}$. Unlike QCNet [50], which generates trajectories for future vehicle motions over a fixed period, our world model adopts an auto-regressive approach, predicting vehicles’ step-wise motion based on its previous predictions. This approach enables the modeling of how past movements influence future decisions.
@@
Figure 11: Illustration of the decoder network module. We have $N$ attention layers to decode accurate trajectories. (1) Agent-to-Map Cross Attention incorporates the map information $z^{m}=\xi(\mathcal{M})$ ($\xi$ is a map encoder) including all kinds of lane markings and types of polygons into the most recent agents’ features such that e.g., $\boldsymbol{z}^{s}_{i,t}=\text{CrossAttn}(\tau_{1,t},\dots,\tau_{{I},t},z^{m})$. (2) Agent-to-Temporal Cross Attention that embeds the historical information into the agents’ trajectories $\tau_{1,t+1},\dots,\tau_{{I},t+1}$ where $\tau_{1,t+1}$ consists of $a_{1,t}$ and $s_{1,t}$, e.g., $z^{{W}}_{i,t}=\text{CrossAttn}(\{z^{s}_{i,t-1},\dots,z^{s}_{i,t-{W}}\},z^{s}_{i,t})$, where $H$ indicates the length of encoding historical steps. (3) Agent-to-Neighbor Cross Attention that embeds the spatial-temporal features of the surrounding agents such that $z^{N}_{i,t}=\text{CrossAttn}(z^{{W}}_{i,t},\{z^{{W}}_{j,t}\}_{j\in N_{i}})$ and (4) Self-Agent Attention considers the agent’s entire predicted trajectory at each time step, capturing dependencies between distant elements, e.g., $z^{A}_{i,t}=\text{SelfAttn}(\{z^{N}_{i,t}\}^{T}_{t=0})$. The final state $s_{t+1}$ is represented by concatenating $s_{t}=\{\boldsymbol{z}^{A}_{i,w}\}_{w=t-{W},i=0}^{t,{I}}$ with the predicted $\boldsymbol{z}^{A}_{1,t+1},\dots,\boldsymbol{z}^{A}_{{I},t+1}$.
@@
Under the context of Dec-POCMDP, when $t=0$, $s_{0}$ captures the static game map $\mathcal{M}$ and initial features of all agents. when ${t>0},s_{t}=\{\boldsymbol{z}^{A}_{i,w}\}_{w=t-{W},i=0}^{t,{I}}$ captures the spatial-temporal information of all agents under the game map $\mathcal{M}$ in the previous $t-1$ time steps, and $a_{1,t},\dots,a_{i,t}$ denotes the acceleration and heading of agents in the current time step. Under this setting, our decoder is implemented by (1) Agent-to-Map Cross Attention, (2) Agent-to-Temporal Cross Attention, (3) Agent-to-Neighbor Cross Attention and (4) Self-Agent Attention, which incorporates map information , historical information , the spatial-temporal features of the surrounding agents , and the features of the agent itself into the temporal dimension thereby mapping $s_{t}$ and $\boldsymbol{a}_{1,\dots,{I},t}$ to $s_{t+1}$. The details of the implementation are illustrated in Figure 11. In addition, our world model also includes two other important modules:
@@
The actor model predicts the actions (acceleration and heading) based on the state $s_{t}=\{\boldsymbol{z}^{s}_{i,w}\}_{w=t-{W},i=0}^{t,{I}}$ to satisfy latent traffic rules in the realistic driving scenarios. The actor model $\pi_{i,t}(a_{i,t}|o_{i,t})$ denotes the probability of that $i$’s agent generates an action $a_{t}$ such that ${\max\limits_{k}}\enspace\omega_{i}^{k}p(a_{t}|\mu_{i,t}^{k},b_{i,t}^{k})=\pi_{i,t}(a_{i,t}|o_{i,t})$ where (1) $\omega_{i}^{k}$ denotes a learnable coefficient and (2) $p_{k}$ denotes the $k$-th mixture component’s Laplace density. We constrain the output actions within a reasonable range to prevent irrational driving behaviors such as sudden acceleration or deceleration, sharp turns, etc. In the subsequent RL fine-tuning stage, we sample the $i$-th agent’s future trajectory as a weighted mixture of Laplace distributions by following [50, 63]:
@@
where $\omega_{i}^{k}$ can effectively act as the weighting coefficients and $\mu_{i,t}^{k}$ and $b_{i,t}^{k}$ characterize the mean position and the level of uncertainty, respectively, of the $i$-th agent at the time step $t$.
@@
For each agent, our observation model maps $s_{t}$ and $a_{i,t}$ onto agent-specific observations $\{\boldsymbol{o}_{i,t}\}_{i}^{I}$. The observation model is implemented by $o_{i,t}=f^{\text{MLP}}_{i}(z^{A}_{i,t})$, where MLP is the Multilayer Perception. Additionally, this module considers the historical actions and observation of all agents such that $\boldsymbol{h}=\{(o_{i,\iota},a_{i,\iota})\}_{\iota=0,i=1}^{w,{I}}$.
@@
As our environment is structured as a multi-player competitive game with rewards and costs specific to each agent, we were inspired by [40, 43] to consider a decentralized update of each agent’s policy where we fix the rest ${I}-1$ agents’ policy $\boldsymbol{\pi}_{-i}$ and train policy $\pi_{i}$ to get the best response of agent $i$. In this paper, we updated $\pi_{i}$ by iteratively optimizing the following objective:
@@
where $\beta_{i}(o)=\frac{c}{\rho_{i}(o)}$ serves as an exploration bonus to less visited state. $c$ is a hyper-parameter, and $\rho_{i}(o)$ denotes the density of visited observation $o$ [64], representing the probability of $o$ occurrences at time $t$ [65]. Such an optimistic V-learning objective serves as an extension to the CCE-V-Learning algorithm [40, 38], which has been proven to converge to CCE under discrete environments, and we extend this algorithm to solve continuous decision-making problems.
@@
We follow [43] and incorporate the Bregman divergence $\mathcal{B}_{\psi}(\cdot,\cdot)$ with respect to the mirror map $\psi$ such that $\mathcal{B}_{\psi}(x,y)=\psi(x)-\psi(y)-\langle\nabla\psi(y),x-y\rangle$ into the objective with convergence guarantees. Recent studies [44, 43, 41, 42] confirmed that the mirror decent approaches can solve different kinds of games in multi-player settings. To derive a more intuitive objective, we implement the mirror map as the negative entropy such that $\psi(x)=\sum p(x)\log p(x)$, and the objective (12) becomes:
@@
However, both the transition function $\mathcal{T}$ and policy of other players $\boldsymbol{\pi}_{-i,t}$ are not subject to optimization in the objective (12), and thus recent studies [66, 67] have often considered the discounted causal entropy [68] $\sum_{t=0}^{T}\gamma^{t}\mathcal{H}[\pi(a_{i,t}|o_{i,t})]$. Similarly, instead of utilizing the computationally intractable trajectory-level KL-divergence $\mathcal{D}_{kl}(\pi_{i}\|\pi^{j-1}_{i})$, we consider the time-wise causal KL-divergence $\sum_{t=0}^{T}\gamma^{t}\mathcal{D}_{kl}[\pi_{i,t}(\cdot)\|\pi^{j-1}_{i,t}(\cdot)]$, and by substituting it and the equation (11) into the objective (12), we have:
@@
To effectively simulate complex traffic scenarios that include various vehicle types, such as cars, buses, and trucks, and diverse driving styles, such as aggressive and conservative driving, it is crucial to tailor the behavior of each agent to regulate the level of competition within the scenarios created. We can more thoroughly assess the system’s robustness by subjecting the AV control system to these various scenarios. This comprehensive evaluation helps in developing trustworthy AV vehicles capable of performing reliably in realistic traffic conditions. To accurately represent the diverse levels of scenarios characterized by different CCEs, we incorporate constrained and risk-sensitive policy optimization into the multi-agent traffic simulation system.
To accurately represent diverse levels of scenarios characterized by the CCEs, we incorporate the constrained and risk-sensitive policy optimization into the multi-agent traffic simulation system.
@@
To dynamically adjust the intensity of CCEs, we request the agents to impose varying levels of driving constraints on the AV agents, thereby modulating the severity and nature of the driving conditions. Specifically, we expand the objective (15) by formulating the trade-off between rewards and costs under a constrained policy optimization objective:
@@
This strategy explicitly manages the risk sensitivity of driving behaviors, thereby deriving risk-seeking or risk-averse policies for each AV agent. This approach effectively promotes either aggressive or conservative driving policies to enhance the realism and variability of the scenarios. With the aim of guiding risk-sensitive and constraints-satisfying policies for multiple agents, inspired by [55], we develop a risk-sensitive constraint to capture the uncertainty induced by environmental dynamics and extend the Objective 16 as follows:
@@
Following [71], we parameterize the distribution with $N$ supporting quantiles and update these function via quantile regression [56], which acts as an asymmetric squared loss in an interval [$-\kappa,\kappa$] around zero: